Solve the system of equations given: x+3z=12, -x-2y+z=10, 3x+5y+2z=-7 ... the answer in came up with is x+y=5-2z.
Answer by Alan3354(69443)

Solve the system of equations given: x+3z=12, -x-2y+z=10, 3x+5y+2z=-7 ... the answer in came up with is x+y=5-2z.
----------
Solve the system means find values of x, y & z.
As in:
x =
y =
z =
----------
x+3z=12
-x-2y+z=10
3x+5y+2z=-7
----------
The 1st eqn has no y term, so eliminate y by combining Eqns 2 & 3.
-x-2y+z=10 times 5 --> -5x - 10y + 5z = 50
3x+5y+2z=-7 times 2 --> 6x + 10y + 4z = -14
-----
-5x - 10y + 5z = 50
+6x + 10y + 4z = -14
------------------------------ Add
x + 9z = 36
x + 3z = 12
--------------- Subtract
6z = 24
z = 4 ********************
x + 3z = 12
x = 0 ****************
y = -3 **************